A(n)-2.0 pC charge is located at (-0.80 m, 0 m) and a(n) 5.0 pC charge is located at (0.40 m, 0 m). Find the net force on a(n) 8.0 uC charge located at (0 m, 1.0 m) 13.magnitude 0.3315 N 0.2935 N 0.2778 N 0.3010 N 0.2812 N 0.3016 N C. 14. direction 127.8° 143.3° 101.6 156.20 116.8° 137.6° C.
Answer #1

F1sine1 01 F1 F1cose F2sine2 62 q1--2 0.8m 0.4m F2cose2

q_{1}=-2mu C=2*10^{-6}C

q_{2}=5mu C=5*10^{-6}C

q_{3}=8mu C=8*10^{-6}C

d_{1}=sqrt{0.8^{2}+1^{2}}

di-1.281m

d) = V 0.42 + 12

d2 = 1.077m

===============

F_{1}=kq_{1}q_{3}/d_{1}^{2}

Fi9 102 10810/(1.281m)-

F_{1}=9*10^{9}*2*10^{-6}*8*10^{-6}/1.281^{2}

F10.144/1.2812

F1 0.08775N

=================

F_{2}=kq_{2}q_{3}/d_{2}^{2}

F2 9105 10810-6C/(1.077m)2

F_{2}=9*10^{9}*5*10^{-6}*8*10^{-6}/1.077^{2}

F_{2}=0.36/1.077^{2}

F, 0.31M

========

F_{x}=-F_{2}cos heta _{2}-F_{1}sin heta _{1}

F_{x}=-0.31*0.4/1.077-0.08775*0.8/1.281

Frー -0.1699M

-------------------------

F_{y}=F_{2}sin heta _{2}-F_{1}cos heta_{1}

F_{y}=0.31*1/1.077-0.08775*1/1.281

{color{Red} F_{y}=0.2193N}

------------

F=sqrt{F_{x}^{2}+F_{y}^{2}}

F=sqrt{(-0.1699)^{2}+(0.2193)^{2}}

F-0.2774N

ANSWER :Option c. 0.2778 N

==========================

Direction

tan heta =F_{y}/F_{x}

heta =tan^{-1}(F_{y}/F_{x})

heta =tan^{-1}(0.2193/-0.1699)

θ=-52 (2nd quadrant)

heta =180^{circ}-52.23^{circ}=127.766^{circ}

ANSWER: Option a.127.8 deg

========================